# J-PLUS: Spectroscopic validation of H emission line maps in spatially resolved galaxies

### v1

@@ 4.1 J-IFU module
The J-IFU module builds the H +[NII] emaps of J-PLUS galaxies. The input of the pipeline are the J-PLUS observation field tile information (tile id and zero points), and galaxy coordinates (RA and Dec). The initial step in the pipeline involves downloading the J-PLUS images of the tile in filters that contain the requested galaxies. This is followed by PSF homegeneisation on all images of the field. This process is done using JYPE-PSF, which is described in the next section. After the PSF homogenisation, we construct the emaps by following the recipe of three filter (3F) method (see Sect. 4.1.3) from Vilella-Rojo et al. (2015). After building the emap of each J-PLUS observation field, we cropped the image to a small size by including the extent of the galaxy for further analysis. The flow chart of the module is given in Fig. 4 (top-left flowchart) and more details are described below.
@@ 4.1.1 PSF homogenization
JYPE-PSF first analyze the full width at half-maximum (FHWM) of the entire tile and select an optimal target PSF as the highest FWHM measure in that tile. Using PSFEx, it derives the PSF homogenization kernel for the input coadded images. Then the convolution kernel K(x) is applied to the PSF model, and minimizes (based on ) the difference with the target PSF for the entire tile. The homogenization kernel components are stored as a FITS data cube. The target shape is a perfectly round analytical function of the Moffat profile. Finally, each coadded image is convolved using Fast Fourier Transform (FFT) with this kernel to homogenize the variable PSF to a constant arbitrary shape to the target PSF.
@@ 4.1.2 Zero point map
In the 3F method (Vilella-Rojo et al., 2015), two broad-band filters (rSDSS and iSDSS in the case of H are used to trace the continuum and one narrow-band filter (H covering J0660) is used to trace the emission line. Then the H + [NII] flux is derived using Eq.(3) from Vilella-Rojo et al. (2015) as shown below,
The values represent the observed average fluxes inside the filters, while and are defined as:
@@ 4.2 IFU module
After constructing emaps, the 2D images from the IFUs are warped to the pixel grid of the J-PLUS image of the galaxy using astwarp from Gnuastro. The warping pixel scale is set by the highest pixel scale size among the IFUs and J-PLUS. The one with the highest pixel scale size will be used to re-scale the other. PHANGS-MUSE (0.2”/pixel), CALIFA (0.5”/pixel) and MaNGA (0.5”/pixel) have higher spatial sampling compared to J-PLUS (0.55”/pixel), thus the pixel scale of the IFUs are changed to J-PLUS size. Afterwards, the image size of J-PLUS is resampled to the same image size of IFUs. This step is necessary for the pixel-to-pixel comparison between the emaps as we do in later sections.
The examples of output 2D emaps from these two modules and their comparison are given from Figures 5 to 7. The figures also show the full extent of the galaxy emap and the cropped images based on the IFU size. The size of the IFU FoV is labelled in the figure.
@@ 5.1 Comparison of H +[NII] emaps of J-PLUS with other IFUs
The spatially resolved H maps trace the star formation at 10 Myr, ionization mechanisms, mass assembly, and other galaxy properties. The precise continuum subtraction is essential for creating emaps that accurately estimate these physical properties of the observed regions. Here we discuss our findings from the comparison of the H maps constructed from J-SHE pipeline with the other IFUs.
@@ 5.1.3 MaNGA emaps
MaNGA provides high-spectral resolution with a larger spectral coverage for individual galaxies but covers a smaller FoV compared to other IFUs studied here. Due to its smaller field of view at the redshift range studied here, most of the MaNGA observations cover the center of galaxies or part of the galaxy regions. Fig. 7 clearly shows the advantage of J-PLUS FoV compared to that of MaNGA. For this case, J-PLUS can provide the spatially resolved H structure of the full extent of the galaxy in a single observation, while MaNGA only covers the central region of this galaxy. We have demonstrated emap comparison of some of the MaNGA galaxies in Fig. 17.
Note that the cases where the absorption present within the filter windows affects the estimation of the emap and the flux will be underestimated.
@@ 5.3 H emission line regions
The H emission line is crucial to identify the star-forming regions and estimate their star formation rates. The SEDs of these spatially resolved star-forming regions help us to investigate the stellar population within galaxies, and their properties like age, metallicity, and luminosity provide insights into the evolution in galaxies. Because J-PLUS is a large-area photometric survey, it provides an efficient way to map spatially resolved star-forming regions in a large sample of galaxies, offering an alternative to spectroscopy, which is more time consuming. In order to validate our method of building J-PLUS SEDs of H emission line regions, we have compared the SEDs with the spectra from IFUs discussed in Section 2. Later we are planning to fit the SEDs of these clumps to study the stellar population properties.
@@ 5.3.2 Ratio of J-PLUS and IFU H+[NII] fluxes
To quantify the goodness of our extraction method, we analyzed the ratio of the J-PLUS and IFU H+[NII] fluxes of detected emission line clumps, noted . IFU emaps created from pseudo J-PLUS NBs (2D method) is used for this analysis.
@@
In the case of MaNGA, a small part of each galaxy is covered and a few star-forming regions are detected. For the complete sample of MaNGA galaxies, we have detected a total of clumps. Again, we discarded the clumps with S/N less than 10 and also removed those detected in J-PLUS and not present in MaNGA. Explanation for those cases are beyond the scope of this paper. A more detailed study of these clumps will be discussed in a forthcoming article. The red histogram in Fig. 14 shows the distribution of , with a Gaussian fit yielding . After normalizing by the r-ratio, the resulting value of value is . We also calculated the ratios of the i and J0660 filters between J-PLUS and the synthetic J-PLUS filters from MaNGA, obtaining fitting values of , and respectively. The difference in continuum subtracted H fluxes is expected due to the different PSFs between J-PLUS (1.1”) and MaNGA(2.5”). The much larger MaNGA PSF FWHM is due to the hexagonal grid of optical fiber configuration (2”-core diameter fibers separated by 0.5”) of the MaNGA survey which introduce systematic errors and underestimations of the flux measurements.
@@ 6 Conclusions
In this study, we developed and validated the J-SHE pipeline for constructing spatially resolved H+[NII] emission maps and extract the SEDs of every emission line region in nearby galaxies (z 0.0165) in J-PLUS DR3 in a very automated way.
The main findings drawn from our study are summarized below.
The wide FoV, spatial resolution, wavelength coverage, and narrow-band filters of J-PLUS are suitable for spatially resolved spectral studies of nearby galaxies.
@@
As a final note, the Javalambre Physics of the Accelerating Universe Astrophysical Survey (J-PAS; Benitez et al., 2014; Dupke et al., 2016) will observe thousands of square degrees in the northern sky and will provide data for thousands of galaxies with significantly improved spatial resolution thanks to its narrow-band filters in the optical. Its continuous set of filters functions like a low-resolution IFU over large areas, enabling a deeper understanding of galaxy formation and evolution on resolved scales.
@@ Acknowledgements.
RPT thanks the Spanish Ministry of Science and Innovation (MCIN/AEI/10.13039/501100011033 y FEDER, Una manera de hacer Europa) with grant No. PID2021-124918NA-C43, the Governments of Spain and Aragón through their general budgets and the Fondo de Inversiones de Teruel. RPT also thanks the European Union - NextGenerationEU through the Recovery and Resilience Facility program Planes Complementarios con las CCAA de Astrofísica y Física de Altas Energías - LA4. RPT acknowledge David Fernández Gil and Francisco Arizo Borillo for their valuable inputs during the analysis of this work. ACS acknowledges funding from the brazilian agencies Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) and the Rio Grande do Sul Research Foundation (FAPERGS) through grants CNPq-314301/2021-6, FAPERGS- 24/2551-0001548-5. AE acknowledges the financial support from the Spanish Ministry of Science and Innovation and the European Union - NextGenerationEU through the Recovery and Resilience Facility project ICTS-MRR-2021-03-CEFCA. AAC acknowledges financial support from the Severo Ochoa grant CEX2021- 001131-S funded by MCIN/AEI/10.13039/501100011033 and the Spanish project PID2023-153123NB-I00, funded by MCIN/AEI. Based on observations made with the JAST80 telescope and T80Cam camera for the J-PLUS project at the Observatorio Astrofísico de Javalambre (OAJ), in Teruel, owned, managed, and operated by the Centro de Estudios de Física del Cosmos de Aragón (CEFCA). We acknowledge the OAJ Data Processing and Archiving Unit (UPAD) for reducing the OAJ data used in this work. Funding for the J-PLUS Project has been provided by the Governments of Spain and Aragón through the Fondo de Inversiones de Teruel; the Aragonese Government through the Research Groups E96, E103, E16_17R, E16_20R, and E16_23R; the Spanish Ministry of Science and Innovation (MCIN/AEI/10.13039/501100011033 y FEDER, Una manera de hacer Europa) with grants PID2021-124918NB-C41, PID2021-124918NB-C42, PID2021-124918NA-C43, and PID2021-124918NB-C44; the Spanish Ministry of Science, Innovation and Universities (MCIU/AEI/FEDER, UE) with grants PGC2018-097585-B-C21 and PGC2018-097585-B-C22; the Spanish Ministry of Economy and Competitiveness (MINECO) under AYA2015-66211-C2-1-P, AYA2015-66211-C2-2, AYA2012-30789, and ICTS-2009-14; and European FEDER funding (FCDD10-4E-867, FCDD13-4E-2685). The Brazilian agencies FINEP, FAPESP, and the National Observatory of Brazil have also contributed to this project. This work was partly done using GNU Astronomy Utilities (Gnuastro, ascl.net/1801.009) version 0.23.11-45775. Work on Gnuastro has been funded by the Japanese Ministry of Education, Culture, Sports, Science, and Technology (MEXT) scholarship and its Grant-in-Aid for Scientific Research (21244012, 24253003), the European Research Council (ERC) advanced grant 339659-MUSICOS, the Spanish Ministry of Economy and Competitiveness (MINECO, grant number AYA2016-76219-P) and the NextGenerationEU grant through the Recovery and Resilience Facility project ICTS-MRR-2021-03-CEFCA. This work was carried out as part of the PHANGS-MUSE collaboration. Based on observations collected at the European Southern Observatory under ESO programmes 1100.B-0651, 095.C-0473, and 094.C-0623 (PHANGS–MUSE; PI Schinnerer), as well as 094.B-0321 (MAGNUM; PI Marconi), 099.B-0242, 0100.B- 0116, 098.B-0551 (MAD; PI Carollo) and 097.B-0640 (TIMER; PI Gadotti). This study uses data provided by the Calar Alto Legacy Integral Field Area (CALIFA) survey (http://califa.caha.es/). Based on observations collected at the Centro Astronómico Hispano Alemán (CAHA) at Calar Alto, operated jointly by the Max-Planck-Institut fűr Astronomie and the Instituto de Astrofísica de Andalucía (CSIC). Funding for the Sloan Digital Sky Survey IV has been provided by the Alfred P. Sloan Foundation, the U.S. Department of Energy Office of Science, and the Participating Institutions. SDSS-IV acknowledges support and resources from the Center for High Performance Computing at the University of Utah. The SDSS website is www.sdss4.org. SDSS-IV is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration including the Brazilian Participation Group, the Carnegie Institution for Science, Carnegie Mellon University, Center for Astrophysics Harvard & Smithsonian, the Chilean Participation Group, the French Participation Group, Instituto de Astrofísica de Canarias, The Johns Hopkins University, Kavli Institute for the Physics and Mathematics of the Universe (IPMU) / University of Tokyo, the Korean Participation Group, Lawrence Berkeley National Laboratory, Leibniz Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Astrophysik (MPA Garching), Max-Planck-Institut für Extraterrestrische Physik (MPE), National Astronomical Observatories of China, New Mexico State University, New York University, University of Notre Dame, Observatário Nacional / MCTI, The Ohio State University, Pennsylvania State University, Shanghai Astronomical Observatory, United Kingdom Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Oxford, University of Portsmouth, University of Utah, University of Virginia, University of Washington, University of Wisconsin, Vanderbilt University, and Yale University.References
@@
- Bacon et al. (1988) Bacon, R., Adam, G., Baranne, A., et al. 1988, in European Southern Observatory Conference and Workshop Proceedings, Vol. 30, Very Large Telescopes and their Instrumentation, Vol. 2, ed. M. H. Ulrich, 1185
@@
- Vázquez-Mata et al. (2022) Vázquez-Mata, J. A., Hernández-Toledo, H. M., Avila-Reese, V., et al. 2022, MNRAS, 512, 2222

### v2

@@ 4.1 J-IFU module
The J-IFU module builds the H +[NII] emaps of J-PLUS galaxies. The input of the pipeline are the J-PLUS observation field tile information (tile id and zero points), and galaxy coordinates (RA and Dec). The initial step in the pipeline involves downloading the J-PLUS images of the tile in filters that contain the requested galaxies. This is followed by PSF homogenization on all images of the field. This process is done using JYPE-PSF, which is described in the next section. After the PSF homogenization, we construct the emaps by following the recipe of three filter (3F) method (see Sect. 4.1.3) from Vilella-Rojo et al. (2015). After building the emap of each J-PLUS observation field, we cropped the image to a small size by including the extent of the galaxy for further analysis. The flow chart of the module is given in Fig. 4 (top-left flowchart) and more details are described below.
@@ 4.1.1 PSF homogenization
JYPE-PSF first analyze the FWHM of the entire tile and select an optimal target PSF as the highest FWHM measure in that tile. Using PSFEx, it derives the PSF homogenization kernel for the input coadded images. Then the convolution kernel K(x) is applied to the PSF model, and minimizes (based on ) the difference with the target PSF for the entire tile. The homogenization kernel components are stored as a FITS data cube. The target shape is a perfectly round analytical function of the Moffat profile. Finally, each coadded image is convolved using Fast Fourier Transform (FFT) with this kernel to homogenize the variable PSF to a constant arbitrary shape to the target PSF.
@@ 4.1.2 Zero point map
In the 3F method (Vilella-Rojo et al., 2015), two broad-band filters (rSDSS and iSDSS in the case of H are used to trace the continuum and one NB filter (H covering J0660) is used to trace the emission line. Then the H + [NII] flux is derived using Eq.(3) from Vilella-Rojo et al. (2015) as shown below:
The values represent the observed average fluxes inside the filters, while and are defined as
@@ 4.2 IFU module
After constructing emaps, the 2D images from the IFUs are warped to the pixel grid of the J-PLUS image of the galaxy using astwarp from Gnuastro. The warping pixel scale is set by the highest pixel scale size among the IFUs and J-PLUS. The one with the highest pixel scale size will be used to rescale the other. PHANGS-MUSE (0.2”/pixel), CALIFA (0.5”/pixel) and MaNGA (0.5”/pixel) have higher spatial sampling compared to J-PLUS (0.55”/pixel), thus the pixel scale of the IFUs are changed to J-PLUS size. Afterwards, the image size of J-PLUS is resampled to the same image size of IFUs. This step is necessary for the pixel-to-pixel comparison between the emaps as we do in later sections.
The examples of output 2D emaps from these two modules and their comparison are given from Figures 5 to 7. The figures also show the full extent of the galaxy emap and the cropped images based on the IFU size. The size of the IFU FoV is labeled in the figure.
@@ 5.1 Comparison of H+[NII] emaps of J-PLUS with other IFUs
The spatially resolved H maps trace the star formation at 10 Myr, ionization mechanisms, mass assembly, and other galaxy properties. The precise continuum subtraction is essential for creating emaps that accurately estimate these physical properties of the observed regions. Here we discuss our findings from the comparison of the H maps constructed from J-SHE pipeline with other IFUs.
@@ 5.1.3 MaNGA emaps
MaNGA provides high-spectral resolution with a larger spectral coverage for individual galaxies but covers a smaller FoV compared to other IFUs studied here. Due to its smaller field of view at the redshift range studied here, most of the MaNGA observations cover the center of galaxies or part of the galaxy regions. Fig. 7 clearly shows the advantage of J-PLUS FoV compared to that of MaNGA. For this case, J-PLUS can provide the spatially resolved H structure of the full extent of the galaxy in a single observation, while MaNGA only covers the central region of this galaxy. We have demonstrated emap comparison of some of the MaNGA galaxies in Fig. 17. We note that the cases where the absorption present within the filter windows affects the estimation of the emap and the flux will be underestimated.
@@ 5.3 H emission line regions
The H emission line is crucial to identify the star-forming regions and estimate their SFRs. The SEDs of these spatially resolved star-forming regions help us to investigate the stellar population within galaxies, and their properties such as age, metallicity, and luminosity provide insights into the evolution in galaxies. Because J-PLUS is a large-area photometric survey, it provides an efficient way to map spatially resolved star-forming regions in a large sample of galaxies, offering an alternative to spectroscopy, which is more time consuming. In order to validate our method of building J-PLUS SEDs of H emission line regions, we have compared the SEDs with the spectra from IFUs discussed in Section 2. Later we are planning to fit the SEDs of these clumps to study the stellar population properties.
@@ 5.3.2 Ratio of J-PLUS and IFU H+[NII] fluxes
To quantify the goodness of our extraction method, we analyzed the ratio of the J-PLUS and IFU H+[NII] fluxes of detected emission line clumps, noted . IFU emaps created from synthetic J-PLUS NBs (2D method) is used for this analysis.
@@
In the case of MaNGA, a small part of each galaxy is covered and a few star-forming regions are detected. For the complete sample of MaNGA galaxies, we have detected a total of clumps. Again, we discarded the clumps with S/N less than 10 and also removed those detected in J-PLUS and not present in MaNGA. Explanation for those cases are beyond the scope of this paper. A more detailed study of these clumps will be discussed in a forthcoming article. The red histogram in Fig. 14 shows the distribution of , with a Gaussian fit yielding . After normalizing by the r-ratio, the resulting value of value is . We also calculated the ratios of the i and J0660 filters between J-PLUS and the synthetic J-PLUS filters from MaNGA, obtaining fitting values of , and respectively. The difference in continuum subtracted H fluxes is expected due to the different PSFs between J-PLUS (1.1”) and MaNGA(2.5”). The much larger MaNGA PSF FWHM is due to the hexagonal grid of optical fiber configuration (2” core diameter fibers separated by 0.5”) of the MaNGA survey which introduce systematic errors and underestimations of the flux measurements.
@@ 6 Conclusions
In this study, we developed and validated the J-SHE pipeline for constructing spatially resolved H+[NII] emission maps and extract the SEDs of every emission line region in nearby galaxies (z 0.0165) in J-PLUS DR3 in a very automated way. The main findings drawn from our study are summarized below.
The wide FoV, spatial resolution, wavelength coverage, and narrow band filters of J-PLUS are suitable for spatially resolved spectral studies of nearby galaxies.
@@
As a final note, the Javalambre Physics of the Accelerating Universe Astrophysical Survey (J-PAS; Benitez et al., 2014; Dupke et al., 2016) will observe thousands of square degrees in the northern sky and will provide data for thousands of galaxies with significantly improved spatial resolution thanks to its NB filters in the optical. Its continuous set of filters functions like a low-resolution IFU over large areas, enabling a deeper understanding of galaxy formation and evolution on resolved scales.
7 Data availability
Table. A1 is only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.
@@ Acknowledgements.
RPT thanks the Spanish Ministry of Science and Innovation (MCIN/AEI/10.13039/501100011033 y FEDER, Una manera de hacer Europa) with grant No. PID2021-124918NA-C43, the Governments of Spain and Aragón through their general budgets and the Fondo de Inversiones de Teruel. RPT also thanks the European Union - NextGenerationEU through the Recovery and Resilience Facility program Planes Complementarios con las CCAA de Astrofísica y Física de Altas Energías - LA4. RPT acknowledge David Fernández Gil and Francisco Arizo Borillo for their valuable inputs during the analysis of this work. ACS acknowledges funding from the brazilian agencies: Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq) and the Rio Grande do Sul Research Foundation (FAPERGS) through grants CNPq-314301/2021-6, FAPERGS- 24/2551-0001548-5. AE acknowledges the financial support from the Spanish Ministry of Science and Innovation and the European Union - NextGenerationEU through the Recovery and Resilience Facility project ICTS-MRR-2021-03-CEFCA. AAC acknowledges financial support from the Severo Ochoa grant CEX2021- 001131-S funded by MCIN/AEI/10.13039/501100011033 and the Spanish project PID2023-153123NB-I00, funded by MCIN/AEI. Based on observations made with the JAST80 telescope and T80Cam camera for the J-PLUS project at the Observatorio Astrofísico de Javalambre (OAJ), in Teruel, owned, managed, and operated by the Centro de Estudios de Física del Cosmos de Aragón (CEFCA). We acknowledge the OAJ Data Processing and Archiving Unit (UPAD) for reducing the OAJ data used in this work. Funding for the J-PLUS Project has been provided by the Governments of Spain and Aragón through the Fondo de Inversiones de Teruel; the Aragonese Government through the Research Groups E96, E103, E16_17R, E16_20R, and E16_23R; the Spanish Ministry of Science and Innovation (MCIN/AEI/10.13039/501100011033 y FEDER, Una manera de hacer Europa) with grants PID2021-124918NB-C41, PID2021-124918NB-C42, PID2021-124918NA-C43, and PID2021-124918NB-C44; the Spanish Ministry of Science, Innovation and Universities (MCIU/AEI/FEDER, UE) with grants PGC2018-097585-B-C21 and PGC2018-097585-B-C22; the Spanish Ministry of Economy and Competitiveness (MINECO) under AYA2015-66211-C2-1-P, AYA2015-66211-C2-2, AYA2012-30789, and ICTS-2009-14; and European FEDER funding (FCDD10-4E-867, FCDD13-4E-2685). The Brazilian agencies FINEP, FAPESP, and the National Observatory of Brazil have also contributed to this project. This work was partly done using GNU Astronomy Utilities (Gnuastro, ascl.net/1801.009) version 0.23.11-45775. Work on Gnuastro has been funded by the Japanese Ministry of Education, Culture, Sports, Science, and Technology (MEXT) scholarship and its Grant-in-Aid for Scientific Research (21244012, 24253003), the European Research Council (ERC) advanced grant 339659-MUSICOS, the Spanish Ministry of Economy and Competitiveness (MINECO, grant number AYA2016-76219-P) and the NextGenerationEU grant through the Recovery and Resilience Facility project ICTS-MRR-2021-03-CEFCA. This work was carried out as part of the PHANGS-MUSE collaboration. Based on observations collected at the European Southern Observatory under ESO programmes 1100.B-0651, 095.C-0473, and 094.C-0623 (PHANGS–MUSE; PI Schinnerer), as well as 094.B-0321 (MAGNUM; PI Marconi), 099.B-0242, 0100.B- 0116, 098.B-0551 (MAD; PI Carollo) and 097.B-0640 (TIMER; PI Gadotti). This study uses data provided by the Calar Alto Legacy Integral Field Area (CALIFA) survey (http://califa.caha.es/). Based on observations collected at the Centro Astronómico Hispano Alemán (CAHA) at Calar Alto, operated jointly by the Max-Planck-Institut fűr Astronomie and the Instituto de Astrofísica de Andalucía (CSIC). Funding for the Sloan Digital Sky Survey IV has been provided by the Alfred P. Sloan Foundation, the U.S. Department of Energy Office of Science, and the Participating Institutions. SDSS-IV acknowledges support and resources from the Center for High Performance Computing at the University of Utah. The SDSS website is www.sdss4.org. SDSS-IV is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration including the Brazilian Participation Group, the Carnegie Institution for Science, Carnegie Mellon University, Center for Astrophysics Harvard & Smithsonian, the Chilean Participation Group, the French Participation Group, Instituto de Astrofísica de Canarias, The Johns Hopkins University, Kavli Institute for the Physics and Mathematics of the Universe (IPMU) / University of Tokyo, the Korean Participation Group, Lawrence Berkeley National Laboratory, Leibniz Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Astrophysik (MPA Garching), Max-Planck-Institut für Extraterrestrische Physik (MPE), National Astronomical Observatories of China, New Mexico State University, New York University, University of Notre Dame, Observatário Nacional / MCTI, The Ohio State University, Pennsylvania State University, Shanghai Astronomical Observatory, United Kingdom Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Oxford, University of Portsmouth, University of Utah, University of Virginia, University of Washington, University of Wisconsin, Vanderbilt University, and Yale University.References
@@
- Bacon et al. (1995) Bacon, R., Adam, G., Baranne, A., et al. 1995, A&AS, 113, 347